An object dropped from rest on a planet takes 7.17 seconds to reach a velocity of -58.1 m/s. The acceleration due to gravity is?


1
Expert's answer
2021-11-21T17:24:46-0500

We can find the acceleration due to gravity on the planet from the kinematic equation:


"v=v_0+gt,""g=\dfrac{v-v_0}{t}=\dfrac{-58.1\ \dfrac{m}{s}-0}{7.17\ s}=-8.1\ \dfrac{m}{s^2}."

The sign minus means that the acceleration of gravity directed downward.